OpenStudy (anonymous):

If you help, I'll give you a medal :D Verify the identity. cot(x- pi/2) = -tan x

OpenStudy (raden):

cot(x- pi/2) = cot(-(pi/2 - x)) = - cot (pi/2 - x) = - tan(x) proof